OpenStudy (anonymous):

Anyone good with calc graphing? Review problem i need some help with: Find the equation of the tangent line to the curve y=1⁄x at the point B(2,1⁄2). Can someone walk me through the steps for this?

OpenStudy (campbell_st):

can you find the 1st derivative..?

OpenStudy (anonymous):

y-y0=m(x-x0). y0=1/2, x0=2, m=y'(2)

OpenStudy (anonymous):

what i had was y= 1/2x- 1/2 but thats wrong based off that i guess...

OpenStudy (campbell_st):

ok... thats not quite correct... if you write the problem in index form its \[y = x^{-1}\] this is what you need to find the derivative of...

OpenStudy (campbell_st):

so apply the differentiation rule \[y = x^n...then...y' = n \times x^{n -1}\]

OpenStudy (anonymous):

I guess I'm having issues with my labeling in these words problems then, can we change the integers and do an example walk through so I can see ?

OpenStudy (anonymous):

Because it makes no sense what I had down as the n value apparently

OpenStudy (campbell_st):

ok... so n = -1 \[y' = -1 \times x^{-1 -1} = -x^{-2}\] does that make sense..?

OpenStudy (campbell_st):

which can be written as \[y' = \frac{-1}{x^2}\]

OpenStudy (anonymous):

that makes sense to me from prior problem right, for the variable labeling how do you know what exactly n is when given the coordinates and y?

OpenStudy (campbell_st):

ok... so n is just the power of x. So what you have from the 1st derivative is the equation for the slope of the tangent. to get the specific slope at the point (2, 1/2) just substitute x = 2 into the 1st derivative. what do you get...?

OpenStudy (anonymous):

@campbell_st You've always been a smart person! I thank you in a large amount for all the help you do on OpenStudy! I just had to say it.

OpenStudy (agent0smith):

" for the variable labeling how do you know what exactly n is when given the coordinates and y?" You're overthinking it. n is the power on x. That's it. Has nothing to do with the points given. When finding the derivative of y=1⁄x ignore EVERYTHING ELSE, all you need is y=1⁄x. No points or anything. The use of the point B(2,1⁄2) comes later. You're mixing too many things together, do one step at a time.

OpenStudy (anonymous):

so I need to solve first what campbell st initially posted?

OpenStudy (anonymous):

y=x^-1 so thats just y= 2^-1 so 1/2 is the 1st?

OpenStudy (agent0smith):

Yep, when finding the derivative, the ONLY thing you should be looking at is \[\Large y = x^{-1}\] then use the power rule

OpenStudy (agent0smith):

"y=x^-1 so thats just y= 2^-1 so 1/2 is the 1st?" where did x go? \[\Large y = x^{-1}\] \[\Large y' = -1 x^{(-1 -1)}\]

OpenStudy (campbell_st):

well not quite \[y' = \frac{-1}{x^2}...x = 2...gives.... y' = \frac{-1}{2^2}\] so the slope of the tangent is m = -1/4 does that make sense?

OpenStudy (campbell_st):

thanks @Mikeyy1992 lots of bright people here who help with maths problems...

OpenStudy (anonymous):

I plugged the 2 into for X which was apparently wrong, thats how i got the 1/2. but i understand now the -1/4 what i did wrong yeah

OpenStudy (anonymous):

Welcome. @campbell_st Yes, I know that very well. It's just amazing to see people put time in helping others.

OpenStudy (anonymous):

i plugged the right value in just wrong format

OpenStudy (campbell_st):

well the slope at the point x = 2 is m = -1/4 so here is a short cut the slope intercept form of a straight line is y = mx + b in your problem you have the tangent equation as y = -1/4 x + b use the point (2, 1/2) and substitute x = 2 and y =1/2 into the tangent equation to find the value of b, and thus the equation of the tangent. hope this helps

OpenStudy (anonymous):

so i had y=-1/4x+ 1 so based off that shortcut im right? since -.5 on the right has to be added to .5 on the left...right?

OpenStudy (campbell_st):

thats the equation I came up with...

OpenStudy (anonymous):

much appreciated

OpenStudy (campbell_st):

glad to help just practice derivatives... and good luck